# Fluid Mechanics Unit Italian Aerospace Research Center (CIRA) Capua (CE), Italy

### v1

@@ Abstract
This paper introduces a family of entropy-conserving finite-difference discretizations for the compressible flow equations. In addition to conserving the primary quantities of mass, momentum, and total energy, the methods also preserve kinetic energy and pressure equilibrium. The schemes are based on finite-difference (FD) representations of the logarithmic mean, establishing and leveraging a broader link between linear and nonlinear two-point averages and FD forms. The schemes are locally conservative due to the summation-by-parts property and therefore admit a local flux form, making them applicable also in finite-volume and finite-element settings. The effectiveness of these schemes is validated through various test cases (1D Sod shock tube, 1D density wave, 2D isentropic vortex, 3D Taylor Green vortex) that demonstrate exact conservation of entropy along with conservation of the primary quantities and preservation of pressure equilibrium.
@@ 1 Introduction
It is generally considered advantageous to discretize the compressible flow equations using supra-conservative [1] schemes that are not only able to conserve the primary invariants, but also additional secondary ones. Aside from the conservation of mass, momentum and total energy, structure-preserving methods for flows typically seek to improve discrete consistency with other properties such as entropy, kinetic energy, pressure-equilibrium, angular momentum, helicity, enstrophy, etc. [2, 3, 4, 5, 6]. In addition to improving the physical relevance and accuracy of the solution field, incorporating more consistencies can improve numerical stability, the benefits of which have been witnessed in increasingly complex physical settings such as for thermally-perfect, multi-component, and reacting gases [7, 8, 9]. For example, kinetic energy serves as a bounded $L_{2}$ estimate on the velocity variables for incompressible flows, while enstrophy serves a similar role with respect to a vorticity variable treatment [10, 11]. Such specialized quantities thus constitute mathematical entropies, which allow for a more rigorous statement on the robustness of the solution—even in the absence of natural stabilization mechanisms such as diffusion. Developing these structure-preserving methods is therefore attractive for achieving a robust yet minimally-dissipative method for use when calculating flows that would otherwise be sensitive to numerical damping, such as in Large Eddy Simulations of high-Reynolds number configurations [12].
@@
For compressible flows, kinetic energy no longer provides a bounded estimate on the flow field. Instead, the thermodynamic entropy ($s$) is often used, as its conservation yields an energy-like statement with respect to a set of special transformation variables (i.e., the entropy variables) [^1] A precise energy statement is available when considering conservation for the class of Harten entropies $h(s)$ [13]. However, unlike the thermodynamic entropy, these functions do not admit provable stability for Navier-Stokes configurations that include Fourier conduction terms.; this notion is furthermore strongly tied to the symmetrization of the Euler equations [14]. To this end, the current work introduces a new class of entropy-conserving methods for the compressible Euler system. Notably, these schemes have the additional property of being compatible with finite differencing.
@@
Entropy-conserving (EC) methods have traditionally been designed in terms of two-point numerical fluxes that are based on a jump condition constraint introduced by Tadmor [15]. Their extension to high-order is then achieved via the notion of flux differencing [16, 17], wherein various low-order two-point fluxes are assembled in a type of Richardson extrapolation. This overall flux framework naturally lends itself to the finite volume (FV) community and has also gained wide use for achieving entropy consistency with modern finite element (FE) schemes [18, 19]. The congruence between flux (i.e., weak) forms and derivative (i.e., strong) forms is generally understood when discretizing linear functions with finite difference (FD) methods [20, 21]. A link between strong and weak forms, however, can also be established for nonlinearly stable discretizations in special instances. For example, the entropy-conservative treatment of Burgers equation can be expressed in a conservative FD form via a quadratic skew-symmetric splitting of the nonlinear term that induces the Heronian mean [22, 23]. While discretely equivalent, the respective methods entail different algorithmic overheads in terms of their implementation, particularly when extending to high-order. Such a direct duality between finite differencing and EC flux forms also exists for the incompressible flow equations [11]; however, the link has not yet been established in the context of the compressible Euler equations and may facilitate the discovery of new schemes.
@@
Although FD formulations for the Euler equations can be designed to preserve linear invariants in addition to other properties like being kinetic-energy-preserving (KEP) and pressure-equilibrium-preserving (PEP), they traditionally do not exhibit the strict EC property. However, a notable class of FD splittings with heightened EC capabilities have been identified, which may be characterized as being quasi entropy-conserving methods (qEC). These include, among others, the recently developed kinetic-energy-entropy-preserving (KEEP) methods [3, 24]. The qEC schemes admit efficient FD representations and have been shown to greatly enhance numerical robustness across compressible regimes despite only being approximately entropy conserving. Interesting to note, is that the flux form associated with these FD-qEC methods emulates the flux form’s structure of Ranocha’s EC scheme [25, 26] [^2] Ranocha’s flux is equivalent to the earlier EC scheme of Chandrashekar [27] in terms of the convective treatment; however, the two methods differ in how the pressure terms are calculated, with Ranocha’s rendition enabling pressure-equilibrium-preservation and also having a direct finite difference representation.; the main point of contrast, however, is how the density ($\rho$) and internal energy ($e$) variables are averaged at the flux interface for the convective flux. A variety of other qEC methods that feature this same structure have recently been proposed [28] and termed asymptotically entropy-conserving; however, these forms are inherently flux-based in nature and are not traditionally associated with a finite difference splitting approach. While FD-qEC methods induce their flux averages (arithmetic or geometric) on the density and internal energy variables via quadratic and cubic splittings of the convective terms [29, 30], this framework does not readily admit the logarithmic-type means that are necessary for recovering strict conservation of the thermodynamic entropy in the case of perfect gases. However, in the same way that standard FD discretizations (with central schemes) of the divergence and advective forms of the convective terms have been linked to numerical fluxes based on bilinear and trilinear interpolations [21, 28], it is of interest to find FD discretizations linked to a Tadmor-type EC method such as Ranocha’s flux.
@@
The current work develops new finite-difference compatible methods that are strictly EC in the context of a single-component ideal gas with the calorically-perfect assumption. In order to achieve this, the intended nonlinear logarithmic flux averages for density and internal energy, $\tilde{\rho}$ and $\tilde{e}$, are first shown to be expressible as quotients of differences and then are explicitly introduced as auxiliary variables in the FD discretization of the compressible Euler system. Rather than satisfying a Tadmor jump condition on the fluxes, the new methods are derived by requiring point-wise cancellation of the spurious volumetric terms that are induced in the discrete entropy equation that is associated with the primary system. Enforcing these cancellations naturally reveals the necessary form of the special auxiliary variables; it also highlights the inherent relation between the choice in FD stencils and the derivative-based representation of the logarithmic means. While it is possible to make FD methods globally entropy conserving via correction procedures [21, 31], the current framework goes beyond by avoiding source terms in the resulting discrete entropy equation. In addition to being EC, the new methods are made to be KEP and PEP as well. Throughout, we employ finite differencing based on diagonal-norm summation-by-parts (SBP) operators [32, 33], including biased renditions [34, 35, 36]. Notably, viable biased EC methods are identified for the first time—a feat that is facilitated by the current FD perspective. A finite difference form of Ranocha’s EC flux is then shown to be recovered as a special symmetric weighting of the two-point forward and backward biased operators.
@@
The paper is organized as follows: Section 2 provides preliminaries in terms of the governing equations and the SBP-based finite difference stencils; Section 3 presents derivative-based representations of specialized nonlinear means; Section 4 introduces the new FD-EC methods; Section 5 confirms the intended scheme properties relative to various test cases (1D density wave, 1D Sod shock tube, 2D isentropic vortex, 3D Taylor-Green vortex); and finally Section 6 provides summary remarks and suggestions for further research.
@@ 2 Preliminaries
$=$ $\dfrac{\partial\rho u_{n}k}{\partial x_{n}}+\dfrac{\partial\rho u_{n}e}{\partial x_{n}}\quad\text{with}\ k=\frac{1}{2}\sum_{m=1}^{d}u_{m}^{2}$
@@ 2.1 Entropy analysis
$ds=c_{v}\frac{dp}{p}-c_{v}\gamma\frac{d\rho}{\rho}\ \to\ s=c_{v}\cdot\log(p/\rho^{\gamma})=c_{v}\cdot\log e-c_{v}(\gamma-1)\cdot\log\rho\ ,$ (6)
@@
[…]
@@
$=$ $s\cdot\frac{\partial\rho}{\partial t}+\rho\cdot\underbrace{\left(\frac{1}{T}\cdot\frac{\partial e}{\partial t}-\frac{p}{\rho^{2}T}\cdot\frac{\partial\rho}{\partial t}\right)}_{\partial s/\partial t}$
@@
We will focus on the spatial discretization of the system of Eqs. (1)–(3) in the context of a semidiscretized procedure. This means that the analysis will assume that temporal integration can be carried out at the continuous level, whereas spatial approximation is obtained by using a discrete method. The framework of a conservative Finite Difference (FD) method is adopted, where variables constitute the nodal values of the continuous unknown functions over a (typically uniform) Cartesian mesh. The formulation used is conservative in the sense that convective and derivative terms are approximated by numerical schemes that can be expressed as sum of differences of numerical fluxes at adjacent nodes along each Cartesian direction. This property discretely reproduces the divergence structure of the convective terms in Eqs. (1)–(7) and is hence referred to as the discretization’s local conservation property. Global conservation (i.e., the property that the total amount of the balanced quantity is not affected by the convective terms, except for boundary contribution) is then guaranteed by the telescoping property, mimicking what happens at the continuous level.
@@ 2.2 Conservative finite difference stencils
[…]
@@
where $(\cdot)_{\ell/r}$ correspond to consistent approximations of the function at the left and right boundaries. Note that on the interior of the domain, away from boundaries, the $\delta^{\text{c}}$ stencils above are central and asymmetric. In the case of classic finite difference SBP operators featuring nodes at the boundary, then $(\cdot)_{\ell/r}$ is the nodal value at the boundary itself such that $u_{\ell}v_{\ell}=(uv)_{\ell}$; however, additional interpretations are possible via the generalized SBP formalism [32] which accommodates the interpretation of such values as being consistent extrapolations or quasi-interpolations of the near boundary data. The above SBP statement can be further extended to biased dual-pair SBP operators with local difference stencils $\delta^{\pm}$ [38, 34, 36] as
@@
with the relationship that
$\delta^{\text{c}}=(\delta^{+}+\delta^{-})/2$ (13)
and where $(\delta^{+}-\delta^{-})$ recovers an artificial dissipation term. In the above, $h_{i}$ are scaling factors associated with a diagonal-norm of the SBP operator and can be interpreted as a quadrature rule [39]. Rearranging terms in Eqs. (11) and (12) reveals the quadratically-split “advection form" (i.e., $u\delta v+v\delta u$) that has been shown to yield a conservative telescoping like representation [20, 40, 37]. The nominal “divergence form" (e.g., $\delta(uv)$) also features a conservative telescoping property. In either case, general finite-volume fluxes can be derived from the SBP-compatible finite difference stencils via a recursive relation
@@
It is important to note that the interior advective form’s flux for the two-point central operator is symmetric in its arguments; however, this is not the case for the biased splitting, which would provide additional flexibility in tailoring the behavior of the discretizations—for example with respect to interacting with sharp gradients in the solution field etc. In special cases (e.g., interior nodes discretized by central stencils) such two-point fluxes can be combined to express the fluxes associated with high-order operators [20, 41] in the fashion of Richardson type extrapolations. A convex combination of the divergence and advective forms is then possible and known to produce an independent averaging of the inputs. For example, when these forms are evenly weighted, we have
$\begin{array}[]{r c l l}\frac{1}{2}\left[\delta_{\text{2pt}}(uv)+u\delta_{\text{2pt}}v+v\delta_{\text{2pt}}u\right]&\longrightarrow&\frac{1}{2}\left[\mathcal{F}(u,v)_{\text{div}}+\mathcal{F}(u,v)_{\text{adv}}\right]_{i+1/2}&=\frac{u_{i+1}+u_{i}}{2}\frac{v_{i+1}+v_{i}}{2}\\ \\ \frac{1}{2}\left[\delta^{+}_{\text{2pt}}(uv)+u\delta^{+}_{\text{2pt}}v+v\delta^{-}_{\text{2pt}}u\right]&\longrightarrow&\frac{1}{2}\left[\mathcal{F}^{+}(u,v)_{\text{div}}+\mathcal{F}^{+}(u,v)_{\text{adv}}\right]_{i+1/2}&=\frac{u_{i+1}+u_{i}}{2}v_{i+1}\\ \\ \frac{1}{2}\left[\delta^{-}_{\text{2pt}}(uv)+u\delta^{-}_{\text{2pt}}v+v\delta^{+}_{\text{2pt}}u\right]&\longrightarrow&\frac{1}{2}\left[\mathcal{F}^{-}(u,v)_{\text{div}}+\mathcal{F}^{-}(u,v)_{\text{adv}}\right]_{i+1/2}&=\frac{u_{i+1}+u_{i}}{2}v_{i}\\ \\ \frac{1}{2}\left[\delta^{+}_{\text{2pt}}(uv)+u\delta^{-}_{\text{2pt}}v+v\delta^{+}_{\text{2pt}}u\right]&\longrightarrow&\frac{1}{2}\left[\mathcal{F}^{+}(u,v)_{\text{div}}+\mathcal{F}^{-}(u,v)_{\text{adv}}\right]_{i+1/2}&=u_{i+1}\frac{v_{i+1}+v_{i}}{2}\\ \\ \frac{1}{2}\left[\delta^{-}_{\text{2pt}}(uv)+u\delta^{+}_{\text{2pt}}v+v\delta^{-}_{\text{2pt}}u\right]&\longrightarrow&\frac{1}{2}\left[\mathcal{F}^{-}(u,v)_{\text{div}}+\mathcal{F}^{+}(u,v)_{\text{adv}}\right]_{i+1/2}&=u_{i}\frac{v_{i}+v_{i+1}}{2}\\ \\ \end{array}$ (37)

### v2

# Italian Aerospace Research Center (CIRA) Capua (CE), Italy

@@ Abstract
This paper introduces a family of entropy-conserving finite-difference discretizations for the compressible flow equations. In addition to conserving the primary quantities of mass, momentum, and total energy, the methods also preserve kinetic energy and pressure equilibrium. The schemes are based on finite-difference (FD) representations of the logarithmic mean, establishing and leveraging a broader link between linear and nonlinear two-point averages and FD forms. The schemes are locally conservative due to the summation-by-parts property and therefore admit a local flux form, making them applicable also in finite-volume and finite-element settings. The effectiveness of these schemes is validated through various test cases (1D Sod shock tube, 1D density wave, 2D isentropic vortex, 3D Taylor–Green vortex) that demonstrate exact conservation of entropy along with conservation of the primary quantities and preservation of pressure equilibrium.
@@ 1 Introduction
It is generally considered advantageous to discretize the compressible flow equations using supra-conservative [1] schemes that are not only able to conserve the primary invariants, but also additional secondary ones. Aside from the primary conservation of mass, momentum and total energy, structure-preserving methods for flows typically seek to improve discrete consistency with other secondary properties such as entropy, kinetic energy, pressure-equilibrium, angular momentum, helicity, enstrophy, etc. [2, 3, 4, 5, 6, 7]. In addition to improving the physical relevance and accuracy of the solution field, incorporating more consistencies can improve numerical stability, the benefits of which have been witnessed in increasingly complex physical settings such as for thermally perfect, multi-component, and reacting gases [8, 9, 10, 11, 12]. For example, kinetic energy serves as a bounded $L_{2}$ estimate on the velocity variables for incompressible flows, while enstrophy serves a similar role with respect to a vorticity variable treatment [13, 14]. Such specialized quantities thus constitute mathematical entropies, which allow for a more rigorous statement on the robustness of the solution—even in the absence of natural stabilization mechanisms such as diffusion. In other instances, methods have been designed to directly satisfy entropy dynamics in order to emphasize nonlinear stability and are then instead made to recover primary conservation of total energy as an auxiliary consequence [15, 16, 17]. Overall, developing such structure-preserving methods is attractive for achieving a robust yet minimally dissipative method for use when calculating flows that would otherwise be sensitive to numerical damping, such as in Large Eddy Simulations of high-Reynolds number configurations [18].
@@
In incompressible flow, kinetic energy acts as a bounded $L_{2}$ measure of velocity, while enstrophy provides similar control in vorticity-based formulations [13, 14]. These quantities function as mathematical entropies and can be used to assess nonlinear stability. For compressible flows, kinetic energy no longer provides a bounded estimate on the flow field. Instead, the thermodynamic entropy ($s$) is often used, as its conservation yields an energy-like statement with respect to a set of special transformation variables (i.e., the entropy variables) [^1] A precise energy statement is available when considering conservation for the class of Harten entropies $h(s)$ [19]. However, unlike the thermodynamic entropy, these functions do not admit provable stability for Navier–Stokes configurations that include Fourier conduction terms.; this notion is furthermore strongly tied to the symmetrization of the Euler equations [20]. To this end, the current work introduces a new class of entropy-conserving methods for the compressible Euler system. Notably, these schemes have the additional property of being compatible with finite differencing.
@@
Entropy-conserving (EC) methods have traditionally been designed in terms of two-point numerical fluxes that are based on a jump condition constraint introduced by Tadmor [21]. Their extension to high-order is then achieved via the notion of flux differencing [22, 23], wherein various low-order two-point fluxes are assembled in a type of Richardson extrapolation. This overall flux framework naturally lends itself to the finite volume (FV) community and has also gained wide use for achieving entropy consistency with modern finite element (FE) schemes [24, 25]. The congruence between flux (i.e., weak) forms and derivative (i.e., strong) forms is generally understood when discretizing linear functions with finite difference (FD) methods [26, 27]. A link between strong and weak forms, however, can also be established for nonlinearly stable discretizations in special instances. For example, the entropy-conservative treatment of Burgers’ equation can be expressed in a conservative FD form via a quadratic skew-symmetric splitting of the nonlinear term that induces the Heronian mean [28, 29]. While discretely equivalent, the respective methods entail different algorithmic overheads in terms of their implementation, particularly when extending to high-order. Such a direct duality between finite differencing and EC flux forms also exists for the incompressible flow equations [14]; however, the link has not yet been established in the context of the compressible Euler equations and may facilitate the discovery of new schemes.
@@
Although FD formulations for the Euler equations can be designed to preserve linear invariants in addition to other properties like being kinetic-energy-preserving (KEP) and pressure-equilibrium-preserving (PEP), they traditionally do not exhibit the strict EC property. However, a notable class of FD splittings with heightened EC capabilities have been identified, which may be characterized as being quasi entropy-conserving methods (qEC). These include, among others, the recently developed kinetic-energy-entropy-preserving (KEEP) methods [3]. The qEC schemes admit efficient FD representations and have been shown to greatly enhance numerical robustness across compressible regimes despite only being approximately entropy conserving. Interesting to note, is that the flux form associated with these FD-qEC methods emulates the flux form’s structure of Ranocha’s EC scheme [30, 31] [^2] Ranocha’s flux is equivalent to the earlier EC scheme of Chandrashekar [32] in terms of the convective treatment; however, the two methods differ in how the pressure terms are calculated, with Ranocha’s rendition enabling pressure-equilibrium-preservation and also having a direct finite difference representation, which yields greater energy consistency with respect to the pressure-velocity coupling.; the main point of contrast, however, is how the density ($\rho$) and internal energy ($e$) variables are averaged at the flux interface for the convective flux. A variety of other qEC methods that feature this same structure have recently been proposed [33, 34, 35] and may be termed asymptotically entropy-conserving; however, these forms are inherently flux-based in nature and are not traditionally associated with a finite difference splitting approach. While FD-qEC methods induce their flux averages (arithmetic or geometric) on the density and internal energy variables via quadratic and cubic splittings of the convective terms [36, 37], this framework does not readily admit the logarithmic-type means that are necessary for recovering strict conservation of the thermodynamic entropy in the case of perfect gases. However, in the same way that standard FD discretizations (with central schemes) of split forms based on the divergence and advective forms of the convective terms have been linked to numerical fluxes based on bilinear and trilinear interpolations [27, 34], it is of interest to find FD discretizations linked to a Tadmor-type EC method such as Ranocha’s flux.
@@
The current work develops new finite-difference compatible methods that are strictly EC in the context of a single-component ideal gas with the calorically perfect assumption. In order to achieve this, the concept of split forms, which is traditionally based on the application of the product rule, is generalized by using the chain rule applied to suitable nonlinear transformations. This is then used to generate different expressions for the discretized convective terms that are equivalent at the continuous limit. In this way, the intended nonlinear logarithmic flux averages for density and internal energy, $\tilde{\rho}$ and $\tilde{e}$, are first shown to be expressible as quotients of differences and then are linked to nonlinear numerical fluxes. Rather than satisfying a Tadmor jump condition on the fluxes, the new methods are derived by requiring point-wise cancellation of the spurious volumetric terms that are induced in the discrete entropy equation that is associated with the primary system. Enforcing these cancellations naturally reveals the necessary form of the special auxiliary variables; it also highlights the inherent relation between the choice in FD stencils and the derivative-based representation of the logarithmic means. While it is possible to make FD methods globally entropy conserving via correction procedures [38, 27, 39, 40, 41, 42, 43] (or cell-wise for element-based methods), the current framework goes beyond by avoiding source terms in the resulting discrete entropy equation. In addition to being EC, the new methods are made to be KEP and PEP as well. Throughout, we employ finite differencing based on diagonal-norm summation-by-parts (SBP) operators [44, 45], including biased renditions [46, 47, 48]. Notably, viable biased EC methods are identified for the first time—a feat that is facilitated by the current FD perspective. A finite difference form of Ranocha’s EC flux is then shown to be recovered as a special symmetric weighting of the two-point forward and backward biased operators.
@@
The paper is organized as follows: Section 2 provides preliminaries in terms of the governing equations and the SBP-based finite difference stencils; Section 3 presents derivative-based representations of specialized nonlinear means; Section  4 introduces the new FD-EC methods; Section 5 confirms the intended scheme properties relative to various test cases (1D density wave, 1D Sod shock tube, 2D isentropic vortex, 3D Taylor–Green vortex); and finally Section 6 provides summary remarks and suggestions for further research. Then in the Appendix, we furthermore supply the following: the entropy conservation proof in matrix-vector notation with SBP operators (Appendix A), the schemes expressed in Hadamard product form (Appendix B), an example of incorporating entropy stable artificial dissipation (Appendix C), and employing the biased two-point fluxes in a multi-point fashion via flux differencing (Appendix D).
@@ 2 Preliminaries
$=$ $\dfrac{\partial\rho u_{n}k}{\partial x_{n}}+\dfrac{\partial\rho u_{n}e}{\partial x_{n}}\quad\text{with}\ k=\frac{1}{2}\sum_{m=1}^{d}u_{m}^{2}.$
@@ 2.1 Entropy analysis
$\mathrm{d}s=\frac{\mathrm{d}e}{T}-R\frac{\mathrm{d}\rho}{\rho}\ \to\ s-s_{\text{ref}}=c_{v}\cdot\log(e/e_{\text{ref}})-c_{v}(\gamma-1)\cdot\log(\rho/\rho_{\text{ref}})\ ,$ (6)
@@
$w\triangleq\frac{\partial\rho s}{\partial\underbrace{[\rho,\rho u_{m},\rho E]}_{[q_{1},\dots,q_{2+d}]}}=\frac{1}{T}\cdot\left[sT-e-\frac{p}{\rho}+\frac{1}{2}\sum_{m=1}^{d}u_{m}^{2},\ -u_{m},\ 1\right]=\underbrace{[w_{\rho},\ w_{\rho u_{m}},\ w_{\rho E}]}_{[w_{1},\dots,w_{2+d}]}$ (8)
@@
$=$ $s\cdot\frac{\partial\rho}{\partial t}+\rho\cdot\underbrace{\left(\frac{1}{T}\cdot\frac{\partial e}{\partial t}-\frac{p}{\rho^{2}T}\cdot\frac{\partial\rho}{\partial t}\right)}_{\partial s/\partial t}.$
@@
We will focus on the spatial discretization of the system of Eqs. (1)–(3) in the context of a semidiscretized procedure. This means that the analysis will assume that temporal integration can be carried out at the continuous level, whereas spatial approximation is obtained by using a discrete method. The framework of a conservative Finite Difference method is adopted, where variables constitute the nodal values of the continuous unknown functions over a (typically uniform) Cartesian mesh. The formulation used is conservative in the sense that convective and derivative terms are approximated by numerical schemes that can be expressed as sum of differences of numerical fluxes at adjacent nodes along each Cartesian direction. This property discretely reproduces the divergence structure of the convective terms in Eqs. (1)–(7) and is hence referred to as the discretization’s local conservation property. Global conservation (i.e., the property that the total amount of the balanced quantity is not affected by the convective terms, except for boundary contribution) is then guaranteed by the telescoping property, mimicking what happens at the continuous level.
@@ 2.2 Conservative finite difference stencils
In the above, we note that the FD stencil coefficients $d_{i,i+k}$ are dependent on the node at which the derivative is being estimated. Herein, we assume that the progression of stencils is arranged such as to satisfy a Summation-by-Parts (SBP) property [45, 44], which is a discrete form of integration-by-parts, $\int_{\Omega}u\,\mathrm{d}v=\left(-\int_{\Omega}v\,\mathrm{d}u+(uv)|_{\partial\Omega}\right)$. Considering a general quadratic decomposition of the input $\phi=uv$, then the SBP property for stencils $\delta^{\text{c}}$ associated with a centered operator is expressed as
@@
where $h_{i}$ are scaling factors associated with a diagonal-norm of the SBP operator, which can be interpreted as a quadrature rule [50], and $(\cdot)_{\ell/r}$ correspond to consistent approximations of the function at the left and right boundaries. In the case of classic finite difference SBP operators featuring nodes at the boundary, then $(\cdot)_{\ell/r}$ is the nodal value at the boundary itself such that $u_{\ell}v_{\ell}=(uv)_{\ell}$; however, additional interpretations are possible via the generalized SBP formalism [44] which accommodates the interpretation of such values as being consistent extrapolations or quasi-interpolations of the near boundary data. Note that in the interior of the domain, away from boundaries, the $\delta^{\text{c}}$ stencils above are central and asymmetric. The above SBP statement can be further extended to biased dual-pair SBP operators with local difference stencils $\delta^{\pm}$ [51, 46, 48] as
@@
with the relationship
$\delta^{\pm}=\overbrace{\frac{1}{2}(\delta^{+}+\delta^{-})}^{\delta^{\text{c}}}\ \pm\ \overbrace{\frac{1}{2}(\delta^{+}-\delta^{-})}^{\delta^{\text{A}D}}\ ,$ (13)
where $\delta^{\text{A}D}$ constitutes an artificial dissipation (AD) term. Rearranging terms in Eqs. (11) and (12) reveals the quadratically split “advection form" (i.e., $u\delta v+v\delta u$) that has been shown to yield a conservative telescoping-like representation [26, 23, 49]. The nominal “divergence form" (e.g., $\delta(uv)$) also features a conservative telescoping property. In either case, general finite-volume fluxes can be derived from the SBP-compatible finite difference stencils via a recursive relation
@@
It is important to note that the interior advective form’s flux for the two-point central operator is symmetric in its arguments; however, this is not the case for the biased splitting, which would provide additional flexibility in tailoring the behavior of the discretizations—for example with respect to interacting with sharp gradients in the solution field etc. In special cases (e.g., interior nodes discretized by central stencils) such two-point fluxes can be combined to express the fluxes associated with high-order operators [26, 52] in the fashion of Richardson-type extrapolations. A convex combination of the divergence and advective forms is then possible and known to produce an independent averaging of the inputs. For example, when these forms are evenly weighted, we have
[…]